OpenStudy (anonymous):

Help please. Will give medal. Choose Always true, sometimes true, or never true. 1) A polynomial divided by a polynomial is a polynomial? (Always true) ?? 2) A polynomial multiplied by a polynomial is a polynomial? 3) A polynomial subtracted by a polynomial is a polynomial?

OpenStudy (igreen):

I believe the 2nd one is always true. Multiplying polynomials will just make the polynomials bigger.

OpenStudy (anonymous):

Ok thanks iGreen do you know if the first one is correct i got Always True?

OpenStudy (igreen):

And I think the first and third one's are sometimes true..

OpenStudy (anonymous):

ok thank you

OpenStudy (igreen):

You could always get 0 by subtracting polynomials. And you can always get a 1 by dividing polynomials. \((2x^2 - 6x + 3) - (2x^2 - 6x + 3) = 0\) \(\dfrac{2x^2 - 6x + 3}{2x^2 - 6x + 3} = 1\)

OpenStudy (igreen):

0 and 1 aren't polynomials.

OpenStudy (anonymous):

Oh ok thank you for explaining. Have a great day.

OpenStudy (anonymous):

@ashhhrodriguez